I am not sure how to answer this...I have never been good at it would you please help me with this problem
A=p(l+rt) solve for t
thank you
pam

Answer by stanbon(75887)

A=p(l+rt) solve for t
---------------
Divide both sides by "p":
1 + rt = A/p
Subtract 1 from both sides:
rt = (A/p) - 1
Divide both sides by "r":
t = [(A/p)-1]/r
=========================
Cheers,
Stan H.